# From pre-transit to post-eclipse: investigating the impact of 3D temperature, chemistry, and dynamics on high-resolution emission spectra of the ultra-hot Jupiter WASP-76b

### v1

@@ 1 Introduction
When it comes to the signature of “3D-ness” in the transmission spectra of UHJs, a relatively coherent picture has emerged. Absorption lines of species that probe the dayside (e.g., metals and CO) tend to show an increasing blueshift during the transit (Ehrenreich et al. 2020; Kesseli et al. 2022; Savel et al. 2023; Wardenier et al. 2023; Pelletier et al. 2023; Prinoth et al. 2023; Beltz et al. 2023; Simonnin et al. 2024). This is a consequence of the strong day-night contrast of the planet in combination with tidally locked rotation. The day-to-night flow produces an additional negative velocity offset of a few km/s. As a result, the detection peaks of these “dayside” species occur at smaller and values compared to the expected planet position in the – map of the observation (Wardenier et al. 2023). For species that mainly probe the nightside of the planet (e.g., H2O), the Doppler shifts of the absorption lines can either become more negative or more positive during the transit (Wardenier et al. 2023, 2024), depending on the atmospheric drag strength. On top of this behavior, asymmetries in scale height (i.e., temperature), cloud cover, or the degree of ionization between the planet’s morning and evening terminator can cause the absorption line strengths to vary substantially during the transit (Wardenier et al. 2021; Savel et al. 2022; Gandhi et al. 2022; Prinoth et al. 2023). Typically, UHJs have a hotter morning limb due to an eastward circulation pattern.
@@ 2.4 Cross-correlation and K–V maps
For each species, we generate two templates, which are derived from the drag-free model: the static dayside spectrum at and the static nightside spectrum at . To obtain these “static” spectra, we set the line-of-sight velocities to zero in the radiative transfer (this means that we also remove the effect of planet rotation from the templates). Because the templates do not depend on orbital phase, they do not account for the fact that different parts of the 3D temperature structure are in view at different phases. Hence, the templates are effectively one-dimensional. Before computing the CCF, we subtract the continuum from both the spectra and the templates by convolving them with a Gaussian kernel and subtracting the result. After this, we “standardize” the individual templates by subtracting the mean and dividing by the standard deviation. We evaluate the CCF for km/s, with steps of 0.25 km/s.
@@
where SNR is the value of the – map at the point , is a scaling factor, and is the number of emission spectra over which the sum is performed. As shown in Fig. 1, for each of the orbital phases that we consider. At each phase angle, we obtain the CCF value at by interpolating between the two values at the nearest velocity shifts in the CCF map.
@@ 3.1 Global circulation models
Fig. 2 shows the temperature structure and the abundance distributions of the three WASP-76b models. The plots depict the 2D planet disk at four different orbital phases and two different pressures ( and bar). In the IGRINS bandpass, the emission spectrum will typically probe pressures between 1 and bar (e.g., Fig. 2 in Smith et al. 2024a). In the optical, Fe lines can probe lower pressures in the millibar regime (e.g., Fig. 4 in Pino et al. 2020). Fig. 3 shows the 1D temperature profiles of the three models in the equatorial plane. As illustrated in both figures, the day-night temperature contrast of the atmosphere increases sharply with altitude. At bar, the dayside temperature is still low enough for H2O to persist. However, at bar, the dayside is completely depleted of H2O due to thermal dissociation. On the other hand, the nightside temperature decreases with altitude, resulting in Fe being condensed out of the atmosphere at bar. Owing to its strong triple bond, CO is neither prone to condensation nor dissociation, leading to a uniform 3D abundance (e.g., Beltz et al. 2022; Savel et al. 2022). Consequently, any changes in the line strength of CO should directly probe the planet’s 3D temperature structure. Finally, Fig. 3 shows that the nightside temperature profile becomes more isothermal with increasing drag strength. As we will demonstrate in the next sections, the strength of the vertical temperature gradient impacts the “detectability” of the nightside spectrum via high-resolution observations.
@@ 3.4 CCF maps of the full orbit
The second column of Fig. 11 shows the CCFs when only accounting for Doppler shifts due to planet rotation. In this case, the net Doppler shifts never exceed the rotational velocity of the planet at the equator, which is indicated by the dashed lines (5.3 km/s for WASP-76b). When the wind profile is taken into account, the Doppler shifts can be stronger. As shown in Fig. 11, the wind profile causes the spectra to be more blueshifted on average.
@@ 3.8 A note on measuring offsets from high-resolution observations
The relatively large uncertainties in the values of some UHJs could be a motivation for more in-depth characterization of their host stars to better constrain their RV semi-amplitude. High-precision spectrographs such as ESPRESSO, MAROON-X, or NIRPS would be ideal for this task. We note, however, that A-type stars such as WASP-33 and WASP-189 are fast rotating and/or active, which could make it significantly harder to lower the error to the km/s level.
Alternatively, a way to circumvent the observational uncertainties in could be to report differences between the values of two species (say, Fe and H2O), thereby “subtracting out” the planet’s orbital velocity.
@@ 3.9 A note on offsets
As shown in Fig. 13, our WASP-76b models do not predict any large offsets in the planet’s dayside emission spectrum, especially when pre- and post-eclipse observations are combined. The reason for this is that the CCF trails in Fig. 11 intersect the line at RV 0 km/s, which is best fit by an orbital trail with km/s. In light of our models, the consistent blueshifts observed in the Fe lines of WASP-76b during pre- and post-eclipse with ESPRESSO ( km/s; Costa Silva et al. 2024) are a bit puzzling. Costa Silva et al. (2024) suggested that this blueshift could be caused by a strong updraft of material around the substellar point of the planet (see their Fig. 8), which effectively moves Fe towards the observer. Our GCMs, however, do not provide strong evidence for such vertical transport, perhaps due to missing physics777One mechanism that our WASP-76b models do not consider is heat transport due to hydrogen dissociation/recombination (e.g., Bell & Cowan 2018; Komacek & Tan 2018; Tan & Komacek 2019; Roth et al. 2021). However, post-processing of the WASP-121b models from Wardenier et al. (2024) (which include hydrogen dissociation/recombination) shows that this additional physics does not lead to stronger offsets in dayside emission spectra. We refer to Appendix B in Bazinet et al. (2025) for the – maps of these models.. Additional forward modeling and follow-up observations with different instruments will be required to resolve the discrepancies between the current data and our models.
@@ 4 Summary and conclusion
J.P.W. acknowledges support from the Trottier Family Foundation via the Trottier Postdoctoral Fellowship, held at the Institute for Research on Exoplanets (IREx). This work was partially funded by the French National Research Agency (ANR) project EXOWINDS (ANR-23-CE31-0001-01). We thank the members of the NIRPS science team for insightful discussions.

### v2

@@ 1 Introduction
When it comes to the signature of “3D-ness” in the transmission spectra of UHJs, a relatively coherent picture has emerged. Absorption lines of species that probe the dayside (e.g., metals and CO) tend to show an increasing blueshift during the transit (Ehrenreich et al. 2020; Kesseli et al. 2022; Savel et al. 2023; Wardenier et al. 2023; Pelletier et al. 2023; Prinoth et al. 2023; Beltz et al. 2023; Simonnin et al. 2024; Seidel et al. 2025). This is a consequence of the strong day-night contrast of the planet in combination with tidally locked rotation. The day-to-night flow produces an additional negative velocity offset of a few km/s. As a result, the detection peaks of these “dayside” species occur at smaller and values compared to the expected planet position in the – map of the observation (Wardenier et al. 2023). For species that mainly probe the nightside of the planet (e.g., H2O), the Doppler shifts of the absorption lines can either become more negative or more positive during the transit (Wardenier et al. 2023, 2024), depending on the atmospheric drag strength. On top of this behavior, asymmetries in scale height (i.e., temperature), cloud cover, or the degree of ionization between the planet’s morning and evening terminator can cause the absorption line strengths to vary substantially during the transit (Wardenier et al. 2021; Savel et al. 2022; Gandhi et al. 2022; Prinoth et al. 2023). Typically, UHJs have a hotter morning limb due to an eastward circulation pattern.
@@ 2.4 Cross-correlation and K–V maps
For each species, we generate two templates, which are derived from the drag-free model: the static dayside spectrum at and the static nightside spectrum at . To obtain these “static” spectra, we set the line-of-sight velocities to zero in the radiative transfer (this means that we also remove the effect of planet rotation from the templates). Because the templates do not depend on orbital phase, they do not account for the fact that different parts of the 3D temperature structure are in view at different phases. Hence, the templates are essentially one-dimensional. Before computing the CCF, we subtract the continuum from both the spectra and the templates by convolving them with a Gaussian kernel and subtracting the result. After this, we “standardize” the individual templates by subtracting the mean and dividing by the standard deviation. We evaluate the CCF for km/s, with steps of 0.25 km/s.
@@
where SNR is the value of the – map at the point , is a scaling factor such that the maximum is 1, and is the number of emission spectra over which the sum is performed. As shown in Fig. 1, for each of the orbital phases that we consider. At each phase angle, we obtain the CCF value at by interpolating between the two values at the nearest velocity shifts in the CCF map.
@@ 3.1 Global circulation models
Fig. 2 shows the temperature structure and the abundance distributions of the three WASP-76b models. The plots depict the 2D planet disk at four different orbital phases and two different pressures ( and bar). In the IGRINS bandpass, the emission spectrum will typically probe pressures as high as bar (e.g., Fig. 2 in Smith et al. 2024a). In the optical, Fe lines can probe lower pressures in the millibar regime (e.g., Fig. 4 in Pino et al. 2020). In Appendix A, we present individual contribution functions for the chemical species considered in this work. Fig. 3 shows the 1D temperature profiles of the three models in the equatorial plane. As illustrated in both Figs. 2 and 3, the day-night temperature contrast of the atmosphere increases sharply with altitude. At bar, the dayside temperature is still low enough for H2O to persist. However, at bar, the dayside is completely depleted of H2O due to thermal dissociation. On the other hand, the nightside temperature decreases with altitude, resulting in Fe being condensed out of the atmosphere at bar. Owing to its strong triple bond, CO is neither prone to condensation nor dissociation, leading to a uniform 3D abundance (e.g., Beltz et al. 2022; Savel et al. 2022). Consequently, any changes in the line strength of CO should directly probe the planet’s 3D temperature structure.
Finally, Fig. 3 shows that the nightside temperature profile becomes more isothermal with increasing drag strength. This can be understood intuitively. Stronger drag leads to less efficient heat redistribution, and thus less energy being deposited onto the nightside. Furthermore, the radiative timescale (the time it takes for a parcel of air to lose a substantial fraction of its energy by radiation) is proportional to , with the photospheric temperature (Perez-Becker & Showman 2013; Parmentier et al. 2021). Because the nightside photosphere is cooler and radiating less energy into space per unit time, radiative equilibrium will converge to a shallower vertical temperature gradient. As we will demonstrate in the next sections, the strength of the vertical temperature gradient impacts the “detectability” of the nightside spectrum via high-resolution observations.
@@ 3.4 CCF maps of the full orbit
The second column of Fig. 11 shows the CCFs when only accounting for Doppler shifts due to planet rotation. In this case, the net Doppler shifts never exceed the rotational velocity of the planet at the equator, which is indicated by the dashed lines (5.3 km/s for WASP-76b). When the wind profile is taken into account, the Doppler shifts can be stronger. As shown in Fig. 11, the wind profile causes the spectra to be more blueshifted on average. In Appendix A, we plot the phase-dependence of the full width at half maximum (FWHM) for all species and all models in pre- and post-eclipse.
@@ 3.8 A note on measuring offsets from high-resolution observations
The relatively large uncertainties in the values of some UHJs could be a motivation for more in-depth characterization of their host stars to better constrain their RV semi-amplitude. High-precision spectrographs such as ESPRESSO, MAROON-X, or NIRPS would be ideal for this task. We note, however, that A-type stars such as WASP-33 and WASP-189 are fast-rotating and/or active, which could make it significantly harder to lower the error to the km/s level.
Alternatively, a way to circumvent the observational uncertainties in could be to report differences between the values of two species (say, Fe and H2O), thereby “subtracting out” the planet’s orbital velocity. However, as shown Fig. 13, differences between the values of two species will typically be small (at most 1-2 km/s in the drag-free model), so these measurements would require high precision to be meaningful.
@@ 3.9 A note on offsets
Our WASP-76b models do not predict any large offsets in the planet’s dayside emission spectrum, especially when pre- and post-eclipse observations are combined. The reason for this is that the CCF trails in Fig. 11 intersect the line at RV 0 km/s, which is best fit by an orbital trail with km/s. In light of our models, the consistent blueshifts observed in the Fe lines of WASP-76b during pre- and post-eclipse with ESPRESSO ( km/s; Costa Silva et al. 2024) are a bit puzzling777Just before submitting the revised version of this manuscript, we learned that a re-analysis of the ESPRESSO data by another team revealed no shifts in , but rather a negative 5 km/s shift in . This result is much more in agreement with our models (G. Guilly, private communication).. Costa Silva et al. (2024) suggested that this blueshift could be caused by a strong updraft of material around the substellar point of the planet (see their Fig. 8), which effectively moves Fe towards the observer. Our GCMs, however, do not provide strong evidence for such vertical transport, perhaps due to missing physics888One mechanism that our WASP-76b models do not consider is heat transport due to hydrogen dissociation/recombination (e.g., Bell & Cowan 2018; Komacek & Tan 2018; Tan & Komacek 2019; Roth et al. 2021). However, post-processing of the WASP-121b models from Wardenier et al. (2024) (which include hydrogen dissociation/recombination) shows that this additional physics does not lead to stronger offsets in dayside emission spectra. We refer to Appendix B in Bazinet et al. (2025) for the – maps of these models.. Additional forward modeling and follow-up observations with different instruments will be required to resolve the discrepancies between the current data and our models.
@@ 4 Summary and conclusion
J.P.W. acknowledges support from the Trottier Family Foundation via the Trottier Postdoctoral Fellowship, held at the Institute for Research on Exoplanets (IREx). We also acknowledge support from the Canadian Space Agency (CSA) under grant 24JWGO3A-03. V.P. is funded by the French National Research Agency (ANR) project EXOWINDS (ANR-23-CE31-0001-01). We thank the members of the NIRPS science team for insightful discussions.
Appendix A Supplementary figures
In this appendix, we present two supplementary figures. Figure 19 shows the rough pressures probed by the emission lines of Fe, CO, H2O, and OH, as derived from the 1D hemispheric dayside average of the drag-free GCM of WASP-76b (see the figure caption for further information about methods). The plots demonstrate that different species probe the atmosphere at a range of pressures between 10-1 and 10-5 bar. While H2O probes relatively deep layers of the atmosphere (pressures higher than 10-3 bar), the Fe emission lines mostly emerge from pressures between 10-3 and 10-5 bar.
Figure 20 shows the full width at half maximum (FWHM) of the CCFs from Fig. 11. During pre-eclipse (), the FWHM increases for all species as a larger part of the dayside rotates into view. That is, the emission signal emerges from an increasingly larger region on the planet disk, probing a wider range of line-of-sight velocities. The opposite effect occurs during post-eclipse (), when parts of the dayside rotate out of view. Comparing the different GCM outputs of WASP-76b, we note that the drag-free model produces the broadest CCF peaks. This is because its strong winds, including the equatorial jet, increase the dispersion in the probed line-of-sight velocities (see Fig. 4). Interestingly, at , the FWHMs of the weak-drag model are 2 km/s lower than the FWHMs of the strong-drag model, suggesting that the day-to-night flow partially counteracts the line-of-sight velocities due to planet rotation. For CO and Fe, the weak-drag model also predicts the CCF peaks to be somewhat broader in post-eclipse than in pre-eclipse.